Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Evaporator_2  Not logged in ELOG logo
Message ID: 687     Entry time: Tuesday, March 25, 2025, 15:26
Author: Darwin Ortiz Rosales 
Target Material: Empty 
Target Oven W.O#: 45447-16 
Source: LP-SIS 
Status: Done 
Subject: TaC coating sintering 

Tuesday, March 25, 2025, 15:25

The coated container as wrapped in 4 Ta foil heatshields.

Vacuum IGP1 initial =e-07 Torr

Target temperature on OMEGA= 1800-1850 C

OMEGA slope:= 0.990

Target time at max temperature = 4 hours

Final Current TBHT= 230 A

Final TGHT current= ~400A (adjustable)

Good vacuum = 5.0 e-04 Torr

Bad vacuum = 5.0E-02 Torr

Will be left overnight to pump down


Wednesday, March 26, 2025, 09:07

starting ramp up at 2A/min to 230A for both TGHT and TBHT


Wednesday, March 26, 2025, 12:48

1800 C at 400A TGHT, 6.9V TGHT

Leaving for 4 hours

ELOG V2.9.2-2455